OpenStudy (anonymous):

Suppose that y depends on x, and x^3-2xy+y^3=1.The derivative dy/dx is given by ?

OpenStudy (anonymous):

3x^2-2dy/dx+3y^2dy/dx=0

OpenStudy (anonymous):

Now solve

OpenStudy (anonymous):

dy/dx=(-3x^2)/(-2+3y^2)

OpenStudy (anonymous):

Thanks!

OpenStudy (zarkon):

\[3x^2-2\left(y+x\frac{dy}{dx}\right)+3y^2\frac{dy}{dx}=0\]

OpenStudy (zarkon):

now solve for dy/dx

OpenStudy (anonymous):

yea got it.Thanks!

OpenStudy (zarkon):

as long as you know the 'answers' above are not correct

OpenStudy (anonymous):

yeaps.i tried.and got the answer in one of the choices given.Thanks a lot btw!!

OpenStudy (zarkon):

ok